A sinusoidal voltage with a period (time

 interval) of 0.1 ms . The radian frequency w

 is:


 select one:

 a. 62830 rad/s

 b. 6283 rad/s

 c. 62.83 rad/s

 POLYTECH

 d. 628.3 rad/s

 e. 6.283 rad/s


1
Expert's answer
2021-06-02T07:43:30-0400

From the problem statement we know

"T=0.1 \cdot 10^{-3}=10^{-4}s"

Then the frequency of radian oscillations is

"w=\frac{2 \cdot \pi}{T}=\frac{2 \cdot 3.1415}{10^{-4}}=62830rad/s"


Answer: a)